# Generator-Based Fuzzers with Type-Based Targeted Mutation

### v2

@@ Abstract
As with any fuzzer, directing Generator-Based Fuzzers (GBF) to reach particular code targets can increase the fuzzer’s effectiveness. In previous work, coverage-guided fuzzers used a mix of static analysis, taint analysis, and constraint-solving approaches to address this problem. However, none of these techniques were particularly crafted for GBF where input generators are used to construct program inputs. The observation is that input generators carry information about the input structure that is naturally present through the typing composition of the program input.
@@
In this paper, we introduce a type-based mutation heuristic, along with constant string lookup, for Java GBF. Our key intuition is that if one can identify which sub-part (types) of the input will likely influence the branching decision, then focusing on mutating the choices of the generators constructing these types is likely to achieve the desired coverages. We used our technique to fuzz AWSLambda applications. Results compared to a baseline GBF tool show an almost 20% average improvement in application coverage, and larger improvements when third-party code is included.
1 Introduction
Fuzz testing (i.e., large-scale automated random testing) can be one of the most practical ways to find bugs that might otherwise be missed by developers.
Coverage-guided fuzzing (hereafter “CGF”), as popularized in the tool AFL [1], augments mutational fuzzing by recording which code locations a test input covers. CGF then prioritizes inputs that cover previously-uncovered code (intuitively, “interesting inputs”) as the starting points for future mutations. CGF is effective at biasing the search process towards relevant inputs without requiring expensive analysis, and this power has made it popular both for practical testing and as the basis for research on further enhancements [2, 3, 4, 5, 6, 7, 8].
@@ 1 Introduction
Coverage-guided fuzzing (hereafter “CGF”), as popularized in AFL [1], augments mutational fuzzing by recording which code locations a test input covers. CGF then prioritizes inputs that cover previously-uncovered code (intuitively, “interesting inputs”) as the starting points for future mutations. CGF is effective at biasing the search process towards relevant inputs without requiring expensive analysis, and this power has made it popular both for practical testing and as the basis for research on further enhancements [2, 3, 4, 5, 6, 7, 8].
However, CGF’s effectiveness is relatively impaired when used to generate structured inputs. This is because traditional CGF lacks information about the input’s structure, leading to the generation of largely syntactically incorrect tests that typically fail to reach the deeper parts of the program’s code. One potential solution to this problem is to use input generators.
@@
An input generator for a given type (also commonly used in property-based testing such as QuickCheck [9]) is just a function that returns a random instance of the type. Using generators to provide inputs yields generator-based fuzzing (hereafter “GBF”). GBF can be used for both unit and system testing, and can capture patterns that constitute legal and typical inputs. The effort of writing generators can be amortized because many data types (e.g., associated with common file formats or libraries) are reused across applications, and generators for composite types can be expressed in terms of generators for their components. More recently, the authors of the Zest system [10] demonstrated how CGF and GBF approaches can be combined by applying mutations to the sequence of random choices a generator uses to build an object. In more detail, generator-based fuzzers work by recording and mutating the sequence of choices made by the pseudo-random number calls within the generator code. We call this sequence of choices the fuzzer chosen input or FCI, as distinct from the outputs of the generator methods which are the inputs to the program under test. Such mutations are more likely to achieve semantically relevant coverage because the generators enforce input validity constraints. However even when CGF and GBF are combined as in Zest, the location for a mutation is still chosen uniformly at random across the input. We observe a missed opportunity that such tools still have no connection between their coverage goals and the process of selecting mutations.
@@
Moreover, more precise targeting of mutations becomes increasingly important as test inputs become larger. On a small input, choosing mutation locations uniformly at random does not require much time to cover every part of an input, just by brute force. But for larger inputs, the most valuable locations to mutate shrink as a fraction of the input size, such as if a format’s metadata has a fixed size while the data it describes gets larger. For large inputs, selecting mutation locations uniformly at random can be wasteful.
While more expensive analysis techniques such as taint tracking or symbolic execution could also be used to trace which parts of the input are relevant to code [11, 12, 13], but they are unlikely to scale. This is because tracking the complete data flow in a cross-language and distributed environment as found in modern applications, i.e., serverless computing, would be very difficult.
@@
Similarly, being selective about relevant code becomes more important when testing larger programs that make extensive use of libraries and frameworks. Libraries and frameworks improve productivity what they encapsulate rich functionality and allow the code that uses them to be relatively simple and high-level. But in this case, the code newly written for an application may be only a small fraction of the total code executed. We will use the term “application” to refer generically to the often smaller code that was written for a specific use case and that is the primary target of fuzz testing. We will use “third-party code” to refer generically to all of the other code that also executes as part of the program under test but is not the primary target of fuzz testing, such as libraries and frameworks. If a fuzzing tool does not differentiate third-party code, then most of the code coverage obligations in CGF will be for the third-party code rather than for the application under test.
@@
To scale fuzz testing as the input size and third-party code usage become significant, fuzzers must concentrate their effort to exercising branches in the application by targeting the most relevant parts of the input.
In this paper, we present a novel technique that can focus the fuzzer to cover application code by mutating substructures in the input that is likely to trigger the intended code target. The core idea of our approach is to use link input generation to code in the application that uses that input via types.
This approach is well-suited for a statically-typed object oriented language like Java, where type information is directly available in the bytecode (or source code) for both input generators and for branches in the program under test. We implement a static analysis that, for each branch in the program under test, determines a ranked list of influencing types of objects that are data-flow predecessors of the branch condition.
As the other half of the connection, our approach associates types with segments of the input, namely the return types of input generator methods. Since generator methods commonly call other generator methods, our implementation uses a simple form of execution indexing [14] to record the type information for input-generating code in the context of its call stack. Our system uses execution indexing to associate a set of generated types with a segment of the fuzzer-controlled input, in a way that maintains that association even if the execution of other generation methods changes.
Our system then uses the matching between types statically associated with branches, and types dynamically associated with parts of the input, to increase the frequency of mutating parts of the input corresponding to objects that influence uncovered branch obligations. As an additional improvement to accelerate the fuzzing process, our static analysis also collects constant strings from the application code, and includes these strings as possibilities when generating strings in program inputs.
@@
For our main evaluation, we test software for serverless microservices written in Java for AWS Lambda and compatible platforms. This is a practically important area where it is common for a small piece of application software to depend on a complex stack of third-party libraries and frameworks. The serverless domain also illustrates why it is appealing to use a lightweight fuzzing technique that does not require entire data flows and software stacks to be instrumented.
@@
The results show that our technique improves the code coverage achieved by fuzzing. Coverage of application branches improved by an average of 18.2% (geometric mean) over the baseline, and coverage including third-party code increased by a geometric mean of 43.2% when the set of covered libraries remained the same, and even more when our approach covers additional libraries. We also separately measure runtime and memory overhead.
@@
The paper is organized as follows: Section 2 provides an example, Section 3 outlines our technique, Section 4 presents our evaluation, Section 5 discusses related work, and Section 6 concludes.
2 Motivating Example
@@
Figure 1: A snippet of a program that creates thumbnails of images for a specific year, written with the help of ChatGPT.
Consider the program in Fig. 1(a), generated with assistance from ChatGPT [15]. The program creates a mosaic image by combining thumbnails of all images modified in the year 2008. Given the location of image files in inputPath (line 7), the program first finds all files within the specified path (line 11), then it iterates over all image files within this directory to collect those whose modification date matches the intended year (lines 11-13). The helper function yearTaken (in Fig. 1(b)) finds the system modification year of a file object. The method first attempts to read attributes of the file system for the image path, particularly the last modified time (line 2-4. Then, it converts that time into a date form captured in a GregorianCalendar object (line 5-6). Finally, the method extracts and returns the year of the last modification of the file (line 7). The thumbnail program in Fig. 1(a) creates a combined thumbnail image for the filtered images (line 15), using some image processing happening within the regroupInSingleImage method, whose details are omitted.
Existing fuzzing tools cannot test this program effectively. This is because this program’s input is not explicitly passed to the program, moreover, the program’s input is substantially larger than inputs to programs typically tested by traditional fuzzers. For example, in this program, multiple images are needed to test the resizing functionality, whereas traditionally fuzzers are used to test a single image processor program. This relatively large input size of the program makes the random mutation strategy unlikely to construct a useful input, i.e., an image with the year 2008, to achieve the desired coverage. In this case, the input bytes representing the date information are likely much smaller than the actual image. For example, to generate an image file of a 100x100 size, where each pixel needs red, blue, and yellow color values, the size of the FCI will be (100x100x4) values. On the other hand, the modification date of the file is about 26 to 30 values within the FCI. This implies the random mutations attempted by GBF are likely to fall outside of the interesting bytes; i.e., mutations are introduced to the actual content of the image, not its file system’s date. Furthermore, observe that the program utilizes some third-party libraries, such as using the ImageIO library for image processing. Also, observe that while all code coverage is potentially relevant and useful for fuzzing, prioritizing coverage of the program under test, in this case the Thumbnail program without the underlying libraries, can potentially be more practical, leading to faster coverage. Thus GBF needs a mechanism to emphasize coverage of coverage targets within the application code while de-emphasizing the significance of third-party library coverage.
2.1 GBF Background
@@
One way to fuzz the program in Fig. 1(a) is by using a generator-based fuzzer, also known as semantic fuzzers [10, 13, 16]. These fuzzers are able to construct structurally correct input using programmable code that we call input generators. For example, Fig. 2 shows a snippet of the code for constructing a file generator Fig. 2(a) FileWithAttributesGenerator which uses, among other generators, the calendar date generator 2(b), i.e.; CalendarGenerator. Both classes define a generate method that constructs the object of the intended type. For example, in the FileWithAttributeGenerator the generate method starts by creating a random file name using the StrGenerator (line 4), then it creates an instance of the FileGenerator, passes the randomly selected name to it, then generates the file’s contents (lines 5-7). Then, the remaining code generates random dates using the CalendarGenerator to set the creation date, last access date, and the last modification dates of the created file. The GregorianCalendar generator creates a calendar object (line 22). It randomly sets its day and month to randomly generated values (lines 23, 24). Finally, it assigns a year to the calendar using a random number that represents a number of years relative to the current year. The selection process employs a geometric distribution with a probability of 1/5. Subsequently, there is a 0.25 probability that the chosen value is added to the current year, resulting in a choosing the current year or a future one, while in other cases, it’s subtracted from the current year, yielding a year in the past.
2.2 Technique Overview
There are two observations on using generators to construct program input. First, observe that the input-generators can naturally be composed together as in the composition of the StrGenerator and CalGenerator within the FileWithAttributesGenerator. Although it is possible to rewrite generators without following this hierarchical arrangement, dividing the object’s generation functionality into smaller, more modular components conforms to sound programming principles and is likely to promote better code organization, reusability, and maintainability. Second, observe that GBF does not directly control the program input. Instead, it controls the random selection within the input-generators represented with operations on r, the SourceOfRandomness. By controlling these random selections, the GBF can indirectly control the generation of program inputs. For example, the GBF fuzzer will control the random selection happening on line 5-7 in Fig. 2(b). GBF can choose between reusing old values or generating fresh random ones, thereby modifying an old input, or constructing a new one.
@@ 2.2 Technique Overview
Making targeted mutations necessitates understanding which parts of the FCI need modification. Directing GBF fuzzers is especially challenging because of the lack of mapping between the FCI and the program’s input. Although heavyweight techniques such as taint analysis, symbolic execution, or certain dataflow analyses can be employed to trace the propagation of the FCI (Fuzzer Chosen Input) to the program’s input, this process is often complex and resource-intensive, potentially impacting the fuzzer’s performance. Additionally, there are scenarios where conducting such analyses may not be feasible due to the presence of heterogeneous languages and systems. For example, it is not easy to track FCI that is going out to a file system.
In this paper, we develop a new type-based mutation heuristic that can direct the mutation to change likely useful subsequences within the FCI based on object types. For example, to cover both arms of the branch at line 12 in Fig. 1(a), one can recognize with a simple static analysis that the object of type GegorianCalendar is the most useful part of the input that needs to be mutated to generate a file whose modification year is 2008. Identifying the types impacting a branch is computed with our influencing type analysis. Then, during the construction of the program input, every byte within the FCI is mapped to the type of objects they are used in generating. For example, suppose $v_{1}=16$ and $v_{2}=8$ are parts of FCI used (lines 23, 24) in Fig. 2(b). Then, one can identify that the dynamic types that $v_{1}$ and $v_{2}$ are part of the GregorianCalendar and File types. This can be computed by looking at the generator’s callstack at the point of generating a random number for these expressions. We call these types the dynamic types. During the fuzzer’s operation, our extension maintains a mapping between the bytes of the FCI and their dynamic types.
During the mutation process, our extension correlates the types influencing the uncovered code targets with the dynamic types linked to the FCI. Subsequently, employing selection heuristics outlined in the subsequent section 3, if a type match is identified, our extension prioritizes mutating the subsequences within the FCI that correspond to the matched type.
@@
We ran the baseline fuzzer and our extension over the program in Fig. 1(a). Each run is an hour long and we repeated the experiment 20 times. Our results show that the baseline fuzzing tool was able to cover line 13 35% of the time (7 out of 20), while our fuzzing extension achieved the same coverage with a success rate of 70% (14 out of 20), with a median 1201.05s (mean 217.9s) versus 998.71s (mean 377s) for the baseline fuzzing tool and our fuzzing extension, showing a skewed probability distribution.
3 Technique

### v3

@@ Abstract
Generator-Based Fuzzers (GBF) is a promising technique that allows fuzzers to construct inputs using programmable generators. This approach can be particularly useful for fuzz applications with complex structured input as it encodes the syntactic structure of the application’s input within its input generators.
However, for a large complex input using a uniform mutation strategy to construct new inputs can have limited effectiveness. This is because, the most valuable locations to mutate within the input shrink as a fraction of its size. This suggests that, as with any fuzzer, directing GBF to reach particular code targets can increase the its effectiveness. In previous work, coverage-guided fuzzers used a mix of static analysis, taint analysis, and/or constraint-solving approaches to address this problem. However, these techniques are unlikely to scale to modern applications. This is because tracking the complete data flow in a cross-language and distributed environment as found in modern applications, e.g., serverless computing, would be very difficult.
@@
In this paper, we introduce a type-based mutation heuristic for Java GBF. Our key intuition is that if one can identify which sub-part (types) of the input will likely influence the branching decision, then focusing on mutating the choices of the generators constructing these types is likely to achieve the desired coverage. We used our technique, along with constant string lookup, to fuzz AWS Lambda applications. Results compared to a baseline GBF tool show an almost 20% average improvement in application coverage, and larger improvements when third-party code is included.
Index Terms:
Testing and Debugging, Program analysis, Testing tools.
I Introduction
Fuzzing (i.e., large-scale automated random testing) can be one of the most practical ways to find bugs that might otherwise be missed by developers, However, even given recent progress, fuzzing is limited in the scale and complexity of modules to which it can effectively be applied. Two techniques, coverage feedback and input generators, help make fuzzing more practical, but we observe a missed opportunity when they are used together for larger inputs and codebases.
@@ I Introduction
Coverage-guided fuzzing (hereafter “CGF”), as popularized in AFL [1], augments mutational fuzzing by recording which code locations a test input covers. CGF then prioritizes inputs that cover previously-uncovered code (intuitively, “interesting inputs”) as the starting points for future mutations. CGF is effective at biasing the search process towards relevant inputs without requiring expensive analysis, and this power has made it popular both for practical bug-finding and as the basis for research on further enhancements [2, 3, 4, 5, 6, 7, 8]. However CGF does not imply any change to the mutation process, which is still random.
@@
Another idea that can improve the effectiveness of fuzzing is to capture patterns of relevant input in code with input generators. An input generator for a given type (also commonly used in property-based testing such as QuickCheck [9]) is just a function that returns a random instance of the type. Using generators to provide inputs yields generator-based fuzzing (hereafter “GBF”). GBF can be used for both unit and system testing, and can capture patterns that constitute legal and typical inputs. The effort of writing generators can be amortized because many data types (e.g., associated with common file formats or libraries) are reused across applications, and generators for composite types can be expressed in terms of generators for their components. More recently, the authors of the Zest system [10] demonstrated how CGF and GBF approaches can be combined by applying mutations to the sequence of random choices a generator uses to build an object. Such mutations are more likely to achieve semantically relevant coverage because the generators enforce input validity constraints. However even when CGF and GBF are combined as in Zest, the location for a mutation is still chosen uniformly at random across the input. We observe a missed opportunity that such tools still have no connection between their coverage goals and the process of selecting mutations.
@@
More precise targeting of mutations becomes increasingly important as test inputs become larger. On a small input, choosing mutation locations uniformly at random does not require much time to cover every part of an input, just by brute force. But for larger inputs, the most valuable locations to mutate shrink as a fraction of the input size, such as if a format’s metadata has a fixed size while the data it describes gets larger. For large inputs, selecting mutation locations uniformly at random can be wasteful.
@@
Similarly, being selective about relevant code becomes more important when fuzzing larger programs that make extensive use of libraries and frameworks. Libraries and frameworks improve productivity what they encapsulate rich functionality and allow the code that uses them to be relatively simple and high-level. But in this case, the code newly written for an application may be only a small fraction of the total code executed. We will use the term “application” to refer generically to the often smaller code that was written for a specific use case and that is the primary target of fuzzing. We will use “third-party code” to refer generically to all of the other code that also executes as part of the program under test but is not the primary target of fuzzing, such as libraries and frameworks. If a fuzzing tool does not distinguish between coverage in the application code and third-party libraries, it may spend most of its time exploring the third-party code. This can lead to an illusion of progress, as increased coverage in the third-party library does not necessarily translate to meaningful coverage in the application code.
@@
For instance, in a program that performs image processing, like the one we present in more detail in Section II, a small amount of code in the application makes decisions based on the image metadata such as its size and creation time. But the metadata is small compared to the image’s pixel data, and the main program is small compared to the third-party image I/O and image processing libraries. To scale fuzzing as file size and third-party code usage become significant, fuzzers must concentrate their effort to exercising branches in the application by targeting the most relevant parts of the input.
@@
For our main evaluation, we fuzz software for serverless microservices written in Java for AWS Lambda and compatible platforms. This is a practically important area where it is common for a small piece of application software to depend on a complex stack of third-party libraries and frameworks. The serverless domain also illustrates why it is appealing to use a lightweight fuzzing technique that does not require entire data flows and software stacks to be instrumented. More expensive analysis techniques such as taint tracking or symbolic execution could also be used to trace which parts of the input are relevant to code [11, 12, 13], but tracking the complete data flow in a cross-language and distributed environment as found in serverless computing would be very difficult.
The core idea of our approach is to link input generation to code in the application that uses that input via types. Part of this link happens statically and another happens dynamically. In the static phase, we use a simple analysis that determine for each branch in the program under test a ranked list of influencing types of objects that are data-flow predecessors of the branch condition. This information is later used to guide the fuzzer toward uncovered application branches, by prioritizing mutating choices within input generators responsible for constructing objects of these statically identified influencing types.
In the dynamic phase, we associate types to dynamically constructed objects from the input generators. Since generator methods commonly call other generator methods to construct complex types, our implementation uses a simple form of execution indexing [14] to record the type information for input-generating code in the context of its call stack.
Finally, to achieve effective mutation, our technique matches types statically associated with uncovered branches, to types dynamically associated with parts of the input. Our technique increases the frequency of mutating parts of the input corresponding to objects of matched influencing types. As in other forms of CGF, this mutation occurs in a loop that measures whether mutated inputs achieve new code coverage, and saves inputs that trigger new coverage to be a basis for future mutations. As an additional improvement to accelerate the fuzzing process, our static analysis also collects constant strings from the application code, and includes these strings as possibilities when generating strings in program inputs.
@@
We evaluated our technique on AWS Lambda applications, as they demonstrate the challenge of fuzzing application code that has a large third-party codebase and loosely defined, but complex and structured, inputs. The results show that our technique improves the code coverage achieved by fuzzing. Coverage of application branches improved by an average of 18.2% (geometric mean) over the baseline, and coverage including third-party code increased by a geometric mean of 43.2% when the set of covered libraries remained the same, and even more when our approach covers additional libraries.
@@
The paper is organized as follows: Section II provides an example, Section III outlines our technique, Section IV presents our evaluation, Section VI discusses related work, and Section VIII concludes.
II Motivating Example
@@
Figure 1: A program snippet that creates thumbnails for images of a specific year.
Consider the program in Fig. 1(a). The program creates a single image composed of thumbnails of all images modified in the year 2008. Given the location of image files in inputPath (line 7), the program first finds all files within the specified path (line 9), then it iterates over all image files within this directory to collect those whose modification date matches the intended year (lines 11-13). The helper function yearTaken (in Fig. 1(b)) finds the system modification year of a file object. The method first attempts to read attributes of the file system for the image path, particularly the last modified time (line 2-4). Then, it converts that time into a date form captured in a GregorianCalendar object (line 5-6). Finally, the method extracts and returns the year of the last modification of the file (line 7). The thumbnail program in Fig. 1(a) creates a combined thumbnail image for the filtered images (line 15), using some image processing happening within the regroupInSingleImage method, whose details are omitted.
Assuming that the fuzzer knows that the input to the program includes all the images in the identified folder, the fuzzer would still have difficulty covering line 13. This is because while the meta-data part of the image files is much smaller than the content of the images and obviously important for achieving coverage of line 13, the fuzzer’s random mutation strategy does not favor modifying the metadata portion of the input.
II-A Generator-Based Fuzzers Background
Figure 2: Generator-Based Fuzzer - FCI is the fuzzer chosen input
To fuzz a program, GBF’s fuzzing driver works as follows (Fig.2). First, the fuzzer driver starts by creating program inputs. This step happens indirectly by running the input generators on the Fuzzer-Chosen Input (FCI), see Fig.2. The FCI captures the random values used by the input generators to construct a random program input. The FCI is consulted whenever the input generator makes a pseudo-random number call. This design allows the GBF to replay previous program inputs, by reusing the same FCI sequences, or create new program inputs by mutating subsequences within the FCI; i.e., using the mutation process. Next, the program under test is executed on the constructed program input, while collecting coverage and output information. Useful inputs, i.e., inputs that achieved new coverages, are saved act as the base of mutation in subsequent fuzzing.
Program inputs generated by this approach are more likely to explore deeper parts of the program under test than those created by traditional CGF. This is because generated program inputs are structurally correct, enabling them to bypass the program’s well-formedness syntax checks. However, because it relies on a uniform random mutation strategy to generate new inputs, finding the right combination to trigger interesting program behavior becomes increasingly difficult as the input space increases.
@@ II-B Fuzzing using GBF
Zest [10] is the state-of-the-art GBF for Java programs. Fig. 3 shows a snippet of the generator’s code written with Zest that constructs a file using a file generator FileWithAttributesGenerator that uses, among other generators, the calendar date generator CalendarGenerator Fig. 3 to generate dates. Both classes define a generate method that constructs the object of the intended type. For example, the FileWithAttributeGenerator’s generate method starts by creating a random filename using the StrGenerator (line 4), then it creates an instance of the FileGenerator that generates the file’s contents (lines 5,6). The remaining code generates random dates (using the CalendarGenerator) to set the creation, the last access, and the last modification dates (lines 9-14). The To create a calendar object, the GregorianCalendar generator choose three random integers for its day, month and year. The year is generated using a geometric distribution with a mean of 5, making it more likely to construct a year that is within $\pm$5 years range from the current year (25%-75% future and past years, respectively) (lines 26-29).
Observe that Zest does not directly control the program input. Instead, it controls, i.e., providing an implementation for, the random selection within the input-generators represented with operations on r, the SourceOfRandomness, highlighted in grey. All such selections is what we refer to as the Fuzzer Controlled Input (FCI). By controlling these random selections, the GBF can indirectly control the generation of program inputs. For example, the GBF fuzzer will control the random selection happening on line 5-7 in Fig. 3.
If one would classify the content of the FCI based on the types they are used in generating, we can see a classification that is similar to the one shown in Fig. 4. The entire FCI is used to generate an object of a File type, which can be further decomposed to subsequences that were used to generate an object of a BufferedImage type (the payload of the image), and the other part is used to generate the meta-data of the file (String and GregorianCalendar types to generate the name and the date information of the file).
The image payload occupies the majority of FIC because it is significantly larger than the metadata. As a result, the fuzzer must make far more random selections to construct the payload than the metadata. For instance, if GBF generates a 100x100 image where each pixel requires red, blue, and yellow color values, the payload segment in FCI would contain 100x100x3 data points. In contrast, the metadata segment would only contain 26-30 data points. This imbalance in the composition of the FCI introduces a bias in mutation, that is, the random uniform mutation strategy will likely introduce modification to image’s payload than its metadata. Consequently, mutations often fall outside the relevant subsequence corresponding to the file system attributes needed to achieve coverage in line 13 in Fig. II. In fact, in a small experiment of an hour long, with 20 repetitions, results show that the baseline FCI was only able to cover line 13, 35% of the time (7 out of 20), while our fuzzing extension achieved the same coverage with a success rate of 70% (14 out of 20).
II-C Fuzzing Using The Proposed Type-Based Mutation
@@
Making targeted mutations necessitates understanding which parts of the FCI need modification. Directing GBF fuzzers is especially challenging because of the lack of mapping between the FCI and the program’s input. Although heavyweight techniques such as taint analysis, symbolic execution, or certain dataflow analyses can be employed to trace the propagation of the FCI within the program, this process is often complex and resource-intensive, potentially impacting the fuzzer’s performance. Additionally, there are scenarios where conducting such analyses may not be feasible due to the presence of heterogeneous languages and systems. For example, it is not easy to track FCI that is going out to a file system.
In this work, we develop a new type-based mutation heuristic that directs mutations toward likely useful subsequences within the FCI, based on the types of objects they are used to generate. Our technique consists of three main steps. First, it statically identifies types that influence branch decisions in the application code. Then, while generating program inputs, it dynamically annotates subsequences within the FCI with the types they contribute to. Finally, it increases the chances of introducing mutations to subsequences whose types influence the decisions of uncovered branches.
To collect influencing types, our technique starts by identifying code targets for the fuzzing process. A code target is defined as the then-side or the else-side of a branching condition within the application code. For example, the then-side and the else-side of the if-statement at line 12 in Fig. 1(a) are two code targets. To cover the then-side of line 12, we statically compute types flowing into a branch. We call this step influencing type analysis which employs a variation of a def-use data flow analysis. This analysis will then associate {GegorianCalendar, File} as the list of influencing types that affects the decision of that code target.
Next, during the construction of the program input, our technique map every value within the FCI to the dynamic type of objects they are used in generating.
Figure 4: FCI used in generating a File from FileWithAttributesGenerator with our extension of Type annotation
For example, Fig. 4 shows the FCI used in generating a File from FileWithAttributesGenerator, annotated by the types they are used in generating. Keeping such a type structure helps us direct the mutation to types that are likely to affect decisions of uncovered code targets. For example, suppose $v_{1}=16$ and $v_{2}=8$ are parts of FCI used (lines 23, 24) in Fig. 3. Then, one can identify that the dynamic types that $v_{1}$ and $v_{2}$ are part of the GregorianCalendar and File types. These type-annotations can be computed by looking at the generator’s callstack at the point of generating a random number for these expressions.
Finally, during mutation, our technique increases the likehood of changing subsequences within the FCI that corresponds to influencing types of uncovered branches. The closer an influencing type is to a code target the more likely it is choose for mutation. For example, our technique would prioritize mutating subsequences used in constructing GregorianCalendar type, since it is the closest influencing type to the code target at line 12.
@@
We ran the baseline fuzzer and our extension over the program in Fig. 1(a). Each run is an hour long and we repeated the experiment 20 times. Our results show that the baseline fuzzing tool was able to cover line 13 35% of the time (7 out of 20), while our fuzzing extension achieved the same coverage with a success rate of 70% (14 out of 20).
III Technique